# Non-Hermitian Multipole Skin Effects Challenge Localization

### v2

@@ Abstract
In their seminal works, Hatano and Nelson studied the effect of disorder on the non-Hermitian skin effect in the non-interacting case Hatano and Nelson (1996); *Hatano1997; *Hatano1998NonHermitian. They found that in open boundary conditions, the system undergoes a transition between two localized phases as a function of disorder strength. At weak disorder, the skin effect persists and wavefunctions are localized at a boundary; at strong disorder, the skin effect is overcome by Anderson localization, and wavefunctions are localized at random positions. When the boundary conditions are periodic, the non-reciprocal hopping creates a persistent current, and the skin effect phase becomes delocalized. In contrast, the Anderson-localized phase remains largely unaffected, since only a few localized modes are impacted by the change in boundary conditions (see Fig. 1(a)). Recently, the stability of this transition has been numerically investigated in interacting systems, which appear to have a non-Hermitian MBL phase Hamazaki et al. (2019); Zhai et al. (2020); Suthar et al. (2022); Ghosh et al. (2022); De Tomasi and Khaymovich (2024); Mák et al. (2024); Roccati et al. (2024).
@@
Here we investigate the impact of disorder on systems that exhibit the non-Hermitian multipole skin effect. First, we generalize the Hatano-Nelson argument to the interacting case and show that, in the presence of U(1) charge conservation, there is a transition between a non-Hermitian skin effect phase and an MBL phase. In the dynamics, this corresponds to a area-to-volume law entanglement phase transition. Next, we consider multipole-conserving systems and demonstrate that, for any amount of disorder, the multipole skin effect remains stable. Remarkably, this implies that under periodic boundary conditions, all eigenstates are delocalized—even in the presence of strong disorder and stringent mobility constraints imposed by multipole conservation.
@@
Figure 2: Non-interacting Hatano-Nelson model, $W=4$. (a) Phase diagram as a function of disorder strength $W$ and non-Hermiticity $g$. (b) An extensive dipole moment in many-body OBC eigenstates indicates the charge skin effect. (c) A finite fraction of PBC energies with a nonzero imaginary part, $f_{imag}$, indicates delocalization (insets: sample single-particle spectra), (d) The imbalance $\mathcal{I}(t)$, after evolving a Néel state in PBC, does not go to zero in the localized phase at long times. Quantities with an overline are averaged over disorder and mid-spectrum eigenstates.
This transition can be understood by making use the similarity transformation Hatano and Nelson (1996)
@@
which maps the non-Hermitian model in OBC to its Hermitian version: $S^{-1}H_{\text{HN}}(g)S=H_{\text{HN}}(0)$. The existence of this transformation automatically implies that the OBC spectrum of $H_{\text{HN}}(g)$ is real. Furthermore, the right eigenstates of the non-Hermitian model can be obtained from the corresponding Hermitian ones: |$\ket{\psi_{g}}=S\ket{\psi_{0}}$. Single-particle states are created by the operators
@@
Now, we turn to the of case of periodic boundary conditions, in which the skin effect phase gives way to delocalization. Heuristically, if the barrier created by the edge is removed, the similarity transformation $S=e^{gP}$ pumps charge preferentially in one direction, leading to a persistent current $j\sim\partial_{t}P$. Without disorder, the energies of delocalized states form a loop in the complex plane: right-movers have a positive imaginary part and are amplified over time, while left-movers have a negative imaginary part and are suppressed. On the other hand, at strong disorder, eigenstates are Anderson localized and therefore cannot be affected by boundary conditions. As a result, the energy of a localized eigenstate under PBC must be the same as its energy under OBC, and hence real. The delocalization transition is therefore associated with a real-to-complex transition of the PBC energy spectrum. The fraction of eigenvalues with a non-zero imaginary part, shown in Fig. 2(c), is therefore a proxy for the transition in PBC. In the Anderson phase, $f_{\text{imag}}\rightarrow 0$, whereas in the delocalized phase, $f_{\text{imag}}\neq 0$.
@@
|$\ket{\psi(t)}=\frac{e^{-iH_{\text{HN}}(g)t}\ket{\psi(0)}}{\lVert e^{-iH_{\text{HN}}(g)t}\ket{\psi(0)}\rVert}.$ (5)
Under PBC, time evolution is frozen in the Anderson localized phase, while a persistent unidirectional current carries charge in the skin effect phase. We contrast these behaviors by starting with a Néel state, |$\ket{\psi(0)}=\prod_{i\in 2\mathbb{Z}}c^{\dagger}_{i}\ket{0}$, and tracking the imbalance of charge over time, ($𝜓𝑡\mathcal{I}(t)=\sum_{i}(-1)^{i}\expectationvalue{n_{i}}{\psi(t)}$. In Fig. 2(d), we show the long time limit of the imbalance, which decays to zero for $g>g_{c}$ since the charges are mobile, but remains finite for $g<g_{c}$.
@@
$\mathcal{O}^{(V=0)}_{\alpha}(0):=$ $\eta^{\dagger}_{\alpha}(0)\eta_{\alpha}(0)=\sum_{i,j}\phi_{\alpha}(i)\phi_{\alpha}^{*}(j)c^{\dagger}_{i}c_{j}$ (7)
$\mathcal{O}^{(V=0)}_{\alpha}(g):=$ $\eta^{\dagger}_{\alpha}(g)\eta_{\alpha}(g)=\sum_{i,j}e^{g(i+j)}\phi_{\alpha}(i)\phi_{\alpha}^{*}(j)c^{\dagger}_{i}c_{j}.$
@@
These operators reduce to Eq. (7) as $V\rightarrow 0$ and fall off exponentially away from the localization center $j_{\alpha}$, such that $\tilde{\phi}_{\alpha}(\{x_{i}\})\sim e^{-\text{max}(|x_{i}-j_{\alpha}|)/\xi}$. All LIOMs commute with the total U(1) charge operator $N=\sum_{j}n_{j}$  [^2] Since both the LIOMs and the charge commute with the Hamiltonian, if $[\mathcal{O}_{\alpha},N]\neq 0$, then there would be a degenerate subspace of states, which is almost impossible due to the onsite disorder. See also Ref. Potter and Vasseur, 2016.. Acting with the similarity transformation, we obtain the corresponding non-Hermitian (right-eigenstate) LIOMs:
@@
In any given term of this expansion, the coefficient of the operator scales as $e^{g\max(x_{i})}e^{-\text{max}(|x_{i}-j_{\alpha}|)/\xi}$, so localization occurs when $g<1/\xi,$ and the skin effect occurs when $g>1/\xi$. Thus, the many-body similarity transformation also explains the existence of non-Hermtian MBL, which was observed in Ref. Hamazaki et al., 2019. We confirm the transition numerically by using the dipole moment as an order parameter (see Fig. 3(a)). When the system is periodic, the skin effect is replaced by a persistent current, and the transition is between MBL and delocalized eigenstates. This transition had been observed numerically in PBC Hamazaki et al. (2019); Mák et al. (2024), though its connection to the skin effect and its explanation in terms of LIOMs were not noted. Note also that a version of the similarity transformation argument was previously used to extract the localization length in a Hermitian MBL system Heußen et al. (2021); O’Brien and Refael (2023).
Interestingly, although the transition connects two localized phases in an OBC system, interactions introduce a crucial difference in the dynamics of quantum entanglement  [^3] We quantify entanglement via the half-partition von Neumann entanglement entropy, $S_{L/2}(t)=-\text{Tr}[\rho_{L/2}(t)\log\rho_{L/2}(t)]$, with $\rho_{L/2}(t)$ the bipartite reduced density matrix of the evolved state $|\psi(t)\rangle$, starting from the Néel state.. In the MBL phase, interaction-induced dephasing drives the system toward a volume-law-entangled state at long times. Conversely, in the skin-effect phase, the state is projected onto its component with maximal dipole moment, resulting in area-law entanglement Gliozzi et al. (2024). The MBL-to-skin-effect transition in OBC therefore induces an entanglement transition from volume- to area-law behavior Kawabata et al. (2023), as shown in Fig. 3(b).
@@
where the kinetic terms describe imbalanced dipole hopping as depicted in Fig. 1(b). We consider the model at half-filling ($N=L/2$) and in the largest dipole sector ($P=0$), and we initially concentrate on the clean limit. Without disorder, this system exhibits a dipole skin effect, whereby dipoles are preferentially pumped to one boundary Gliozzi et al. (2024). Unlike the conventional (charge) skin effect, which clusters charges at one boundary, the dipole skin effect pushes charges symmetrically outwards to both edges of an open chain. In a system without boundaries, this leads to a persistent dipole current and delocalized states (schematically shown in Fig. 1(b)).
@@
One consequence of our results is that multipole skin effects can be stabilized even when multipole conservation is explicitly broken by charge hopping: there exists a regime where disorder localizes the charge while multipoles remain delocalized. Relatedly, it would be interesting to investigate possible transitions between skin effects for different multipoles, like dipoles and charges. The effects of discrete and non-Abelian symmetries on the disordered skin effect should also be explored. Finally, moving beyond one-dimensional systems remains an important avenue for further study.
@@
For U(1) charge conservation, we can “break” the similarity transformation by adding a next-nearest neighbor hopping term to the Hatano-Nelson model of Eq. (1):
@@
In the dipole-conserving case, we can “break” the similarity transformation by choosing different imbalances $g_{0}$ and $g_{1}$ for the range-four and range-five dipole hopping terms in Eq. (12):
@@
In general, our predictions for localization transitions or lack thereof should also be valid for systems in which the similarity transformation does not work, as long the direction of non-reciprocity remains the same. Fundamentally, all multipole skin effects are driven by an imbalanced hopping of $m$-poles. Independently of whether a similarity transformation to a Hermitian Hamiltonian exists, such hopping creates a current of $m$-poles that tends to maximize the $(m+1)$-pole moment in an open system.
@@
Lastly, we confirm that the Hermitian MBL phase at $W=5.5$ has a set of exponentially localized conserved operators. These operators correspond to the LIOMs of Eq. 8 in the main text, and their localization in the Hermitian phase is a key property of MBL. We construct these LIOMS numerically by following Ref. Chandran et al. (2015) and taking the long-time average of a local operator:
@@
We show the behavior of $M_{ij}$ in both the MBL phase and the clean limit in Fig. S3. The results are averaged over 32 disorder realizations and over different sites in a periodic chain of length $L=18$. For a LIOM defined using the full energy eigenbasis as in Eq. (S5), $M_{ij}\rightarrow 0$ at large $|i-j|$ Chandran et al. (2015). However, we construct the $\tilde{n}_{i}$ by only summing over the eigenstates in the symmetry sector of the Néel state. In other words, we actually calculate
@@
Let us restrict our attention to the symmetry sector at half-filling $N=L/2$ and dipole moment $P=0$. Inside this sector, the limited range of the dipole hopping imposes a strong kinetic constraint: many states are not connected to each other by the Hamiltonian. In fact, the sector itself fragments into exponentially many disconnected Krylov subsectors, and the largest of these is exponentially small in system size ($D_{\text{largest}}/D_{\text{total}}\sim e^{-L}$ala et al. (2020); Moudgalya et al. (2020). While in the main text we avoided this (strong) Hilbert space fragmentation by adding a longer-ranged dipole hopping term, it is also interesting to consider how the skin effect and disorder are manifest in individual Krylov subsectors.
The largest Krylov subsector is integrable and can be mapped to a free-fermion chain of length $L/2$ Moudgalya et al. (2020). The mapping involves grouping sites $12k-$ 1 and $2k$ of the original lattice into a new composite site. After this rewriting, the integrable subsector is the one where each pair of sites contains only a single occupied state. In other words, only the computational basis states |:=|$\ket{\rightarrow}:=\ket{01}$ and |:=|$\ket{\leftarrow}:=\ket{10}$ are allowed on every (odd-even) pair of sites. These two states represent a local dipole and a local anti-dipole, respectively. For a composite site $k:=(2k-1,2k)$, we can define dipole creation and annihilation operators, $d^{\dagger}_{k}:=c_{2k-1}c^{\dagger}_{2k}$ and $d_{k}:=c^{\dagger}_{2k-1}c_{2k}$. These operators transform a local dipole into a local anti-dipole and vice versa. When acting on the integrable subspace, they obey fermionic anticommutation relations, so they create genuine local quasiparticle excitations. Furthemore, if we define a number operator for these local dipoles, $n^{d}_{k}=d^{\dagger}_{k}d_{k}$, then the total dipole operator in the subsector can be written as
@@
Such a transition can only occur if the dipoles are local quasiparticles as in this Krylov subsector. Generally, dipoles are 1D objects and their localization (via the dipole skin effect) dominates over the localization of 0D charge in a possible Anderson/MBL phase. However, if dipoles can be mapped to local charges, then, like charged particles, their Anderson localization competes with the skin effect. Nevertheless, this is a highly fine-tuned situation. For the dipole moment to decompose into a sum of local densities as in Eq. (S10), the Hilbert space must be truncated to forbid states containing longer dipoles, and the Hamiltonian must be carefully tuned to preserve these states. For example, even $H_{4}(g)$ does not exhibit this behavior in general, and only a vanishingly small subsector of the Hilbert space hosts a transition.
The hidden U(1) symmetry manifests in another way: the dipole skin effect is more weakly localized than usual inside the Krylov subsector. The dipole skin effect maximizes the quadrupole moment, which generically scales as $L^{3}$ for many-body states, leading to the super-exponential localization of LIOMs discussed in the main text. In this restricted sector, on the other hand, any quadrupole must be formed by local dipoles, and the largest quadrupole moment, associated with the state |=|$\ket{\leftarrow\leftarrow\ldots\rightarrow\rightarrow}=\ket{1010\ldots 0101}$, only scales as $L^{2}$. This weaker scaling implies only exponential localization of LIOMs from the skin effect, and thus allows a localization transition.

### v3

@@ Abstract
In their seminal works, Hatano and Nelson studied the effect of disorder on the non-Hermitian skin effect in the non-interacting case Hatano and Nelson (1996); *Hatano1997; *Hatano1998NonHermitian. They found that in open boundary conditions, the system undergoes a transition between two localized phases as a function of disorder strength. At weak disorder, the skin effect persists and wavefunctions are localized at a boundary; at strong disorder, the skin effect is overcome by Anderson localization, and wavefunctions are localized at random positions. When the boundary conditions are periodic, the non-reciprocal hopping creates a persistent current, and the skin effect phase becomes delocalized. In contrast, the Anderson-localized phase remains largely unaffected, since only a few localized modes are impacted by the change in boundary conditions (see Fig. 1(a)). Recently, the stability of this transition has been numerically investigated in interacting systems, which appear to have a non-Hermitian MBL phase Hamazaki et al. (2019); Zhai et al. (2020); Suthar et al. (2022); Ghosh et al. (2022); De Tomasi and Khaymovich (2024); Mák et al. (2024); Roccati et al. (2024). However, an analytical argument for this MBL phase is lacking, and the fate of intrinsically interacting systems, like those with multipole symmetry, remains unclear.
@@
Here we investigate the impact of disorder on systems that exhibit the non-Hermitian multipole skin effect. First, we generalize the Hatano-Nelson argument to the interacting case and show that, in the presence of U(1) charge conservation, there is a transition between a non-Hermitian skin effect phase and an MBL phase. In the dynamics, this corresponds to an area-to-volume law entanglement phase transition. Next, we consider multipole-conserving systems and demonstrate that, for any amount of disorder, the multipole skin effect remains stable. Remarkably, this implies that under periodic boundary conditions, all eigenstates are delocalized—even in the presence of strong disorder and stringent mobility constraints imposed by multipole conservation.
@@
Figure 2: Non-interacting Hatano-Nelson model, $W=4$. (a) Phase diagram as a function of disorder strength $W$ and non-Hermiticity $g$. Dotted and dashed lines show the scaling of the transition at weak and strong disorder, respectively. (b) An extensive dipole moment in many-body OBC eigenstates indicates the charge skin effect. (c) A finite fraction of PBC energies with a nonzero imaginary part, $f_{imag}$, indicates delocalization (insets: sample single-particle spectra), (d) The imbalance $\mathcal{I}(t)$, after evolving a Néel state in PBC, does not go to zero in the localized phase at long times. Quantities with an overline are averaged over disorder and mid-spectrum eigenstates.
This transition can be understood by making use of the similarity transformation Hatano and Nelson (1996)
@@
[…]
@@
We now turn to the case of periodic boundary conditions, in which the skin effect phase gives way to delocalization. Heuristically, if the barrier created by the edge is removed, the similarity transformation $S=e^{gP}$ pumps charge preferentially in one direction, leading to a persistent current $j\sim\partial_{t}P$. Without disorder, the energies of delocalized states form a loop in the complex plane: right-movers have a positive imaginary part and are amplified over time, while left-movers have a negative imaginary part and are suppressed. On the other hand, at strong disorder, eigenstates are Anderson localized and therefore cannot be affected by boundary conditions. As a result, the energy of a localized eigenstate under PBC must be the same as its energy under OBC, and hence real. The delocalization transition is therefore associated with a real-to-complex transition of the PBC energy spectrum. The fraction of eigenvalues with a non-zero imaginary part, shown in Fig. 2(c), is therefore a proxy for the transition in PBC. In the Anderson phase, $f_{\text{imag}}\rightarrow 0$, whereas in the delocalized phase, $f_{\text{imag}}\neq 0$.
@@
[…]
@@
$\mathcal{O}^{(V=0)}_{\alpha}(0)$= $\eta^{\dagger}_{\alpha}(0)\eta_{\alpha}(0)=\sum_{i,j}\phi_{\alpha}(i)\phi_{\alpha}^{*}(j)c^{\dagger}_{i}c_{j}$ (7)
$\mathcal{O}^{(V=0)}_{\alpha}(g)$= $\eta^{\dagger}_{\alpha}(g)\eta_{\alpha}(g)=\sum_{i,j}e^{g(i+j)}\phi_{\alpha}(i)\phi_{\alpha}^{*}(j)c^{\dagger}_{i}c_{j}.$
@@
These operators reduce to Eq. (7) as $V\rightarrow 0$ and fall off exponentially away from the localization center $j_{\alpha}$, such that $\tilde{\phi}_{\alpha}(\{x_{i}\})\sim e^{-\text{max}(|x_{i}-j_{\alpha}|)/\xi}$. All LIOMs commute with the total U(1) charge operator $N=\sum_{j}n_{j}$  [^2] Since both the LIOMs and the charge commute with the Hamiltonian, if $\mathcal{O}_{\alpha},N]\neq 0$, then there would be a degenerate subspace of states, which is almost impossible due to the onsite disorder. See also Ref. Potter and Vasseur, 2016.. Acting with the similarity transformation, we obtain the corresponding non-Hermitian (right-eigenstate) LIOMs:
@@
In any given term of this expansion, the coefficient of the operator scales as $e^{g\max(x_{i})}e^{-\text{max}(|x_{i}-j_{\alpha}|)/\xi}$, so localization occurs when $g<1/\xi,$ and the skin effect occurs when $g>1/\xi$. Thus, the many-body similarity transformation also explains the existence of non-Hermitian MBL, which was observed in Ref. Hamazaki et al., 2019. We confirm the transition numerically via exact diagonalization, using the dipole moment as an order parameter (see Fig. 3(a)). When the system is periodic, the skin effect is replaced by a persistent current, and the transition is between MBL and delocalized eigenstates. This transition had been observed numerically in PBC Hamazaki et al. (2019); Mák et al. (2024), though its connection to the skin effect and its explanation in terms of LIOMs were not noted  [^3] Other works that numerically explore spectral and dynamical signatures of the non-Hermitian MBL transition include Refs. Zhai et al. (2020); Suthar et al. (2022); Ghosh et al. (2022); De Tomasi and Khaymovich (2024); Roccati et al. (2024), though none discuss the physical mechanism for localization via non-Hermitian LIOMs.. Note also that a version of the similarity transformation argument was previously used to extract the localization length in a Hermitian MBL system Heußen et al. (2021); O’Brien and Refael (2023).
Interestingly, although the transition connects two localized phases in an OBC system, interactions introduce a crucial difference in the dynamics of quantum entanglement  [^4] We quantify entanglement via the half-partition von Neumann entanglement entropy, $S_{L/2}(t)=-\text{Tr}[\rho_{L/2}(t)\log\rho_{L/2}(t)]$, with $\rho_{L/2}(t)$ the bipartite reduced density matrix of the evolved state $|\psi(t)\rangle$, starting from the Néel state.. In the MBL phase, interaction-induced dephasing drives the system toward a volume-law-entangled state at long times. Conversely, in the skin-effect phase, the state is projected onto its component with maximal dipole moment, resulting in area-law entanglement Gliozzi et al. (2024). The MBL-to-skin-effect transition in OBC therefore induces an entanglement transition from volume- to area-law behavior, as shown in Fig. 3(b). Because this entanglement transition relies on disorder, it is distinct from the one previously observed in clean non-Hermitian systems Kawabata et al. (2023).
@@
where the kinetic terms describe imbalanced dipole hopping as depicted in Fig. 1(b). We consider the model at half-filling ($N=L/2$) and in the largest sector of fixed dipole moment ($P=0$), and we initially concentrate on the clean limit. Without disorder, this system exhibits a dipole skin effect, whereby dipoles are preferentially pumped to one boundary Gliozzi et al. (2024). Unlike the conventional (charge) skin effect, which clusters charges at one boundary, the dipole skin effect pushes charges symmetrically outwards to both edges of an open chain. In a system without boundaries, this leads to a persistent dipole current and delocalized states (schematically shown in Fig. 1(b)).
@@
One consequence of our results is that multipole skin effects can be stabilized even when multipole conservation is explicitly broken by charge hopping: there exists a regime where disorder localizes the charge while multipoles remain delocalized. This stability is important for experimental realizations because dipole-conservation is typically only an approximate symmetry. For example, optical lattices with an external tilt effectively conserve dipole moment for long timescales Guardado-Sanchez et al. (2020); Scherg et al. (2021); Kohlert et al. (2023). Onsite disorder is easily implemented in such cold atom systems, and the addition of external loss can induce non-reciprocal couplings Gong et al. (2018), including multipole-hopping Hu et al. (2025). Metamaterials, such as topoelectric circuits Lee et al. (2018) and acoustic lattices Jing et al. (2024), provide another candidate platform, and allow for strong non-reciprocities of up to $g\sim 1.5$.
Relatedly, it would be interesting to investigate possible transitions between skin effects for different multipoles, like dipoles and charges. The effects of discrete and non-Abelian symmetries on the disordered skin effect should also be explored. Finally, moving beyond one-dimensional systems remains an important avenue for further study.
@@
For U(1) charge conservation, we can “break” the similarity transformation by adding a next-nearest neighbor hopping term to the Hatano-Nelson model:
@@
In the dipole-conserving case, we can “break” the similarity transformation by choosing different imbalances $g_{0}$ and $g_{1}$ for range-four and range-five dipole hopping terms:
@@
In general, our predictions for localization transitions or lack thereof should also be valid for systems in which the similarity transformation does not work, as long the direction of non-reciprocity remains the same. Fundamentally, all multipole skin effects are driven by an imbalanced hopping of $m$-poles. Independently of whether a similarity transformation exists, such hopping creates a current of $m$-poles that tends to maximize the $(m+1)$-pole moment in an open system.
@@
Lastly, we confirm that the Hermitian MBL phase at $W=5.5$ has a set of exponentially localized conserved operators. These operators correspond to the LIOMs of Eq. (8) in the main text, and their localization in the Hermitian phase is a key property of MBL. We construct these LIOMS numerically by following Ref. Chandran et al. (2015) and taking the long-time average of a local operator:
@@
We show the behavior of $M_{ij}$ in both the MBL phase and the clean limit in Fig. S3. The results are averaged over 32 disorder realizations and over different sites in a periodic chain of length $L=18$. For a LIOM defined using the full energy eigenbasis as in Eq. (S5), $M_{ij}\rightarrow 0$ at large |ij||i-j| Chandran et al. (2015). However, we construct the $\tilde{n}_{i}$ by only summing over the eigenstates in the symmetry sector of the Néel state. In other words, we actually calculate
@@
Let us restrict our attention to the symmetry sector at half-filling $N=L/2$ and dipole moment $P=0$. Inside this sector, the limited range of the dipole hopping imposes a strong kinetic constraint: many states are not connected to each other by the Hamiltonian. In fact, the sector itself fragments into exponentially many disconnected Krylov subsectors, and the largest of these is exponentially small in system size ($D_{\text{largest}}/D_{\text{total}}\sim e^{-L}$Sala et al. (2020); Moudgalya et al. (2020). While in the main text we avoided this (strong) Hilbert space fragmentation by adding a longer-ranged dipole hopping term, it is also interesting to consider how the skin effect and disorder are manifest in individual Krylov subsectors.
[…]
@@
Such a transition can only occur if the dipoles are local quasiparticles as in this Krylov subsector. Generally, dipoles are 1D objects and their localization (via the dipole skin effect) dominates over the localization of 0D charge in a possible Anderson/MBL phase. However, if dipoles can be mapped to local charges, then, like charged particles, their Anderson localization competes with the skin effect. This is a highly fine-tuned situation. For the dipole moment to decompose into a sum of local densities as in Eq. (S10), the Hilbert space must be truncated to forbid states containing longer dipoles, and the Hamiltonian must be carefully tuned to preserve these states. For example, even $H_{4}(g)$ does not exhibit this behavior in general, and only a vanishingly small subsector of the Hilbert space hosts a transition.
[…]